# Sound Field Synthesis with Acoustic Waves  [^1] Author is affiliated with Amazon Inc. USA.  [^2] This manuscript is an expanded version of a conference paper of the same title

### v1

@@ Abstract
We propose a practical framework to synthesize broadband sound-field on a small rigid surface based on the physics of sound propagation. The sound-field is generated as a composite map of two components: room component and device component; with acoustic plane waves as core element. This decoupling of room and device components significantly reduces the problem complexity and provide accurate rendering of the sound-field. We describe in details the theoretical foundations, efficient procedures, and engineering applications of the proposed framework. The effectiveness of the proposed framework is established through rigorous validation under different environment setups.
@@
Room Acoustics, Sparse Recovery, Acoustic Wave Decomposition.
@@ I Introduction
Sound-field synthesis at a microphone array in a room is the process of synthesizing audio at each microphone of the array from a source signal emanating from a sound source elsewhere in the room. It is a key task in evaluating performance metrics of speech/audio communication devices, as it is a cost-effective methodology for data generation to replace real data collection, which is usually a slow, expensive, and error-prone procedure. Acoustic modeling techniques are frequently utilized to generate synthetic data to either replace or augment real data collection at a fraction of data collection cost. These techniques usually aim at estimating the Room Impulse Response (RIR) between two points in a room. The RIR is either computed empirically using direct measurement, or simulated using a model for room acoustics. Empirical methods are in general accurate, but they are relatively expensive because of the required human labor. The whole procedure for RIR calculation has to be repeated for every source/receiver position combination and for every new device with a different form factor. Moreover, empirical methods require the hardware availability of the device under test which is not always possible especially during hardware development.
@@
Simulation methods address the above issues by relying on computational acoustics rather than measurements to estimate the RIR. A brute-force simulation would solve the inhomogeneous acoustic wave equation with proper boundary conditions of the room and device surface [1]. Though theoretically viable, it requires significant effort to characterize all boundary conditions in a typical room. Further, the simulation time can be prohibitive if it is evaluated over a broadband spectrum. Moreover, the whole simulation needs to be repeated for every new form factor of the device under test. To address the computational complexity, the image source method [2] has been widely used to approximate point-to-point room acoustics. It utilizes the ray tracing concept [1] to significantly reduce modeling and computational complexity of brute-force simulation. Though simple and effective in some scenarios, the image source method has few limitations. For example, it has poor approximation at low frequencies, and it cannot model small surfaces, e.g., furniture, and rough surfaces, e.g., curtains.
@@
In this work, we describe a novel procedure that combines empirical and simulation methods to provide a balanced tradeoff between the two approaches for sound-field synthesis. It splits the sound-field into two independent components: room component, and device component; and the overall sound-field is the composite mapping of the two components. The room component abstracts the room impact at an interior point, due to a predefined sound source in the room, as a superposition of acoustic plane waves; which is computed using a single measurement with a large microphone array. The device component is computed using acoustic simulation or anechoic measurements to evaluate the fingerprint of each acoustic plane wave on the device surface (as measured at the microphone array of the device). The overall acoustic pressure on the device surface when placed at an interior point in a room is computed by plugging in the computed device fingerprints into the acoustic plane wave representation at that point. This arrangement abstracts room acoustics and allows reusing room information with different devices when tested in the same room. Therefore, it enables the concept of room database, which contains abstract room acoustics information that is independent of the device under test. Likewise, it allows reusing the same device component with different rooms. To enable the proposed method, we develop a general procedure to compute the plane wave decomposition at a point in a room by applying sparse recovery techniques on an audio capture with a large microphone array. We also introduce the device dictionary concept, that captures the device acoustic behavior; which is a general concept for general microphone array processing systems [3]. The proposed methodology is rigorously validated across many rooms and many devices with different form factors and microphone array configuration. The synthetic RIR is shown to match the true RIR, in the least square sense, over a broadband spectrum up to $8$ kHz. The synthesis methodology is also shown to closely resemble real measurements in evaluating higher level metrics, e.g., word error rate, and false rejection rate.
@@
The acoustic plane wave expansion has been used in earlier work with model-based sound-field reconstruction, e.g., [4, 5, 6, 7]; where acoustic plane waves are used as kernels for sound-field reconstruction. The plane-wave expansion is interpolated with free-field propagation model to reproduce the sound field within a convex source-free zone. The plane-wave expansion is computed from measurements of an array of microphones placed at the zone perimeter. The computation of the expansion is done either through spherical harmonics or using sparse recovery techniques. In this work, we study a different problem of reproducing the sound field on a rigid surface that is placed at the same point in the room. The key contribution of the current work is introducing the device dictionary concept; which captures the impact of acoustic plane waves on the device surface. This enables the generalization of the sound-field production to a rigid surface with an arbitrary form factor.
@@
The paper is organized as follows. In section II, we lay down the theoretical foundation of the work. The details of the proposed framework are described in section IV. Then, we present the validation results in section V. Finally, we describe few engineering applications in section VI. The following notations are used throughout the paper. A bold lower-case letter denotes a column vector, while a bold upper-case letter denotes a matrix. $M$ always refers to the number of microphones. The independent variables $t$ and $\omega$ refer to time and frequency respectively. Additional notations are introduced when needed.
@@ II-A Acoustic Plane Waves
Acoustic plane waves are eigenfunctions to the homogenous Helmholtz equation, and they constitute a powerful tool for analyzing the wave equation. Further, plane waves are a good approximation of the wave-field emanating from a far-field point source [8]. The acoustic pressure of a plane-wave with vector wave number $\bf{k}$ is defined at a point ${\bf{r}}=(x,y,z)$ in the three dimensional space as:
$\psi({\bf{k}})\triangleq p_{0}e^{-j{\bf{k}}^{T}{\bf{r}}}$ (1)
where ${\bf{k}}$ is the three-dimensional wavenumber vector.
The plane-wave decomposition (PWD) has been used for approximating point-source seismic recording [9, 10, 11], and sound field reproduction [12, 13, 14, 8]. A local solution to the homogenous Helmholtz equation can be approximated by a linear superposition of plane waves of different angles of the form [15, 10]:
$p(\omega)=\sum_{l\in\Lambda}\alpha_{l}\ \psi\left({\bf{k}}(\omega,\theta_{l},\phi_{l})\right)$ (2)
where $\Lambda$ is a set of indices that defines the directions of plane waves $\{\theta_{l},\phi_{l}\}$, each $\psi({\bf{k}})$ is a plane wave as in (1), and $\{\alpha_{l}\}$ are complex scaling factors. We will refer to the wave-field in (2) as the overall background acoustic pressure. The decision variables in this approximation are $\left\{\Lambda,\{\alpha_{l}\}_{l\in\Lambda}\right\}$.
III Device Acoustic Dictionary
Generalizing the free-field plane-wave expansion in (2) to include the scattering due to the device surface, requires computing the device acoustic response to each plane wave. The device response to all plane-wave in the three-dimensional space is collectively referred to as the device acoustic dictionary. The total wave-field at any point on the device surface when the device is impinged by an incident plane-wave $p_{i}({\bf{k}})$ has the general form:
$p_{t}=p_{i}+p_{s}$ (3)
where $p_{t}$ and $p_{s}$ refer to the total and scattered wave-field respectively. $p_{t}$ can be computed numerically by inserting (3) in the Helmholtz equation and solving for $p_{s}$ with appropriate boundary conditions. If a microphone array of size $M$ is mounted on the device surface, and the microphone port size is much smaller than the wavelength, then each microphone can be approximated by a point on the device surface. In this case, the total field, ${\mathbf{p}}_{t}$, at the microphone array, due to an incident plane wave $p_{i}({\mathbf{k}})$, is a vector of size $M$ whose entries are the corresponding total field at each individual microphone. The device acoustic dictionary of a device is composed of vectors of total acoustic wave-field.
@@ III Device Acoustic Dictionary
The output of the above model is the plane-wave dictionary of the device
@@
where each entry in the dictionary is a vector of length $M$, and each element in the vector is the total acoustic pressure at one microphone in the microphone array when a plane wave with ${\bf{k}}(\omega,\theta_{l},\phi_{l})$ hits the device. The dictionary also covers all frequencies of interest typically up to $8$ kHz.
IV Proposed Framework
IV-A Overview
The key idea in the proposed sound synthesis methodology is to generalize the plane-wave expansion in (2), which summarizes room acoustics at a point in a room, to include the impact of scattering due to the device surface. This generalization would yield the combined effect of the room and the device under test, which is the final objective. If the device dimensions are much smaller than the room dimensions, then secondary reflections due to the device surface are negligible; and the device impact on the room acoustics could be ignored. Hence, even after introducing the device into the room, the directions and weights of the acoustic plane waves in (2) do not change. However, because of the device surface, each plane wave $\psi\left({\bf{k}}(\omega,\theta_{l},\phi_{l})\right)$ in (2), has an acoustic fingerprint at the device microphone array; which is the corresponding entry in the device dictionary, $\boldsymbol{\beta}_{l}(\omega)$. Hence, if the device is placed at a point in the room whose sound-field is expressed as in (2), then the observed sound field vector at the device microphone array is
@@ IV-A Overview
The transition from the sound-field in (2) in the absence of the device to the sound field in (5) in the presence of the device is the technical foundation of the proposed synthesis framework. This transition is enabled by the linearity of the wave equation and the introduction of the device acoustic dictionary as described in section III.
Note that, if another device with acoustic dictionary $\mathcal{D}^{(2)}\triangleq\{\boldsymbol{\beta}_{l}^{(2)}(\omega):\forall\ \omega,l\}$ is placed at the same point, then the observed sound-field at the second microphone array
@@
where only the mapping through device dictionary changes, while the directions, $\Lambda$, and weights, $\{\alpha_{l}\}$, of constituent plane waves do not change. This is the essence of the proposed methodology that separates room and device components.
Hence, the three steps for sound-field synthesis at a microphone array mounted on a device that is placed at a point in the room are as follows:
@@
    Compute the free-field plane wave expansion at a point as in (2). This summarizes room acoustics at a point in the room. A procedure for computing this expansion is described in section IV-B. This process is repeated for every new point in a room, and it is independent of the device under test.
@@
    Generate the acoustic dictionary of the device under test as described in section III. This is computed once per device and it is independent of the room.
@@
    For each room position, combine the plane-wave expansion with the device dictionary as in (5) to synthesize the sound-field at the device microphone array.
Repeating step $1$ above for multiple rooms and multiple positions within each room generates a room database; which is generated only once, but could be reused in evaluating and generating data for new devices.
IV-B Acoustic Plane Wave Decomposition
The main technical hurdle in the proposed framework is computing the plane-wave expansion (2) at a point in a room with a source signal emanating from another point in the room. In the proposed framework, this is computed via data capture with a large microphone array of $32$ microphones mounted on a sphere (EigenMike) [20]. The large microphone array is necessary to avoid having an underdetermined system of equations in recovering the constituent plane waves. It was found experimentally that $20$ to $30$ plane waves are sufficient for an accurate approximation of the sound field (with reconstruction error less than $-30$ dB). The plane wave decomposition is composed as an optimization problem whose objective is minimizing the difference in the least square sense between observed and synthesized sound fields. If the observed sound field of the EigenMike at frequency $\omega$, is ${\mathbf{y}}(\omega)$, then the objective function to be minimized has the form
@@ IV-B Acoustic Plane Wave Decomposition
where $\left\{\bar{\boldsymbol{\beta}}_{k}(\omega)\right\}$ are the entries of the EigenMike acoustic dictionary at frequency $\omega$. The decision variables are the set of indices $\Lambda$, and the corresponding weights $\{\alpha_{l}\}$. The L1-regularization in (7) is added to stimulate a sparse solution as $|\Lambda|<30$ is much smaller than the dictionary size, which is in the order of $10^{3}$. The objective function can be put in matrix form as:
@@
IV-C Room Impulse Response (RIR) Computation
RIR aims at modeling the acoustic channel between source and receiver as a linear time-invariant system. The RIR combines both room acoustics and scattering due to device surface, and it is computed once for a given device and a given source/receiver positions in a room. It is a multichannel transfer function where the number of channels equals the size of the microphone array. For RIR computation, a special source signal that covers the whole frequency spectrum, e.g., white noise or Golay sequence [26], is utilized. A source signal with short duration of $2$-$3$ seconds is sufficient for RIR computation. For a source signal $x(t,\omega)$, the EigenMike is utilized to generate the time-frequency map of the plane wave decomposition as described in the previous section. For a device under test, this time-frequency map is combined with the device dictionary to generate the multichannel output signal ${\mathbf{y}}(t,\omega)$ as in (5). The transfer function between $x(t,\omega)$ and ${\mathbf{y}}(t,\omega)$ is computed using system identification techniques. For example, by applying Wiener-Hopf equation in the frequency domain [27], we get

### v2

@@ Abstract
We propose a practical framework to synthesize the broadband sound-field on a small rigid surface based on the physics of sound propagation. The sound-field is generated as a composite map of two components: the room component and the device component, with acoustic plane waves as the core tool for the generation. This decoupling of room and device components significantly reduces the problem complexity and provides accurate rendering of the sound-field. We describe in detail the theoretical foundations, and efficient procedures of the implementation. The effectiveness of the proposed framework is established through rigorous validation under different environment setups.
@@
Room acoustics, acoustic simulation, plane wave decomposition, multichannel audio synthesis.
@@ I Introduction
ound-field synthesis at a microphone array in a room is the process of synthesizing audio at each microphone of the array from a source signal emanating from a sound source elsewhere in the room. It is a key task in evaluating performance metrics of speech/audio communication devices, as it is a cost-effective methodology for data generation to replace real data collection, which is usually a slow, expensive, and error-prone procedure. Acoustic modeling techniques are usually utilized to generate synthetic data to either replace or augment real data collection at a fraction of the cost. These techniques usually aim at estimating the Room Impulse Response (RIR) between two points in the room. The RIR is either computed empirically using direct measurement, or simulated using a model for room acoustics. Empirical methods are in general accurate, but they are relatively expensive because of the required human labor.
@@
Simulation methods provide a cost effective alternative as they utilize computational acoustics rather than physical measurements. A brute-force simulation would solve the inhomogeneous acoustic wave equation with proper boundary conditions of the room and device surface [1]. Though theoretically viable, it requires significant effort to characterize all boundary conditions in a typical room. Further, simulation time can be prohibitive if it is evaluated over a broadband spectrum. Moreover, the whole simulation needs to be repeated for every new form factor of the device under test. To address the computational complexity, the image source method [2] has been widely used to approximate point-to-point room acoustics. It utilizes the ray tracing concept [1] to significantly reduce the modeling and computational complexity of brute-force simulation. Though simple and effective in some scenarios, the image source method has few limitations. For example, it has poor approximation at low frequencies, and it cannot model small surfaces (as compared to wavelength), e.g., furniture, and rough surfaces, e.g., curtains.
@@
In this work, we describe a novel procedure that combines empirical and simulation methods to provide a balanced tradeoff between the two approaches for sound-field synthesis. It splits the sound-field into two independent components: room component, and device component, such that the overall sound-field is the composite mapping of the two components. The room component captures the room impact at an interior point, due to a predefined sound source in the room. This is represented as a superposition of acoustic plane waves, which is computed using a single measurement with a large microphone array. The device component is computed using acoustic simulation or anechoic measurements to evaluate the fingerprint of each acoustic plane wave on the device surface (as measured at the microphone array mounted on the device surface). The overall acoustic pressure on the device surface when placed at an interior point in a room is computed by plugging in the computed device fingerprints into the acoustic plane wave representation at that point. This arrangement provides an efficient representation of room acoustics that allows reusing room information with devices of different form factor when tested in the same room. Therefore, it enables the concept of room database, which contains abstract room acoustics information that is independent of the device under test. Likewise, it allows reusing the same device component with different rooms. To enable the proposed method, we develop a general procedure to compute the plane wave decomposition at a point in a room by applying sparse recovery techniques on an audio capture with a large microphone array. We also utilize the device dictionary concept, that captures the acoustic behavior of general microphone array mounted on a rigid surface of arbitrary form factor [3]. The proposed methodology is rigorously validated across many rooms and many devices with different form factors and microphone array geometry. The synthetic RIR is shown to match the true RIR, in the least square sense, over a broadband spectrum up to $8$ kHz. The synthesis methodology is also shown to closely resemble real measurements in evaluating higher level metrics, e.g., word error rate, and false rejection rate.
@@
The acoustic plane wave expansion has been used in earlier work with model-based sound-field reconstruction, e.g., [4, 5, 6, 7], where acoustic plane waves are used as kernels for sound-field reconstruction. The plane wave expansion is interpolated with free-field propagation model to reproduce the sound field within a convex source-free zone. The plane wave expansion is computed from measurements of an array of microphones placed at the zone perimeter. The computation of the expansion is done either through spherical harmonics or using sparse recovery techniques. In this work, we study a different problem of reproducing the sound field on a rigid surface that is placed at the same point in the room. A key contribution of the current work is utilizing the device dictionary concept for sound synthesis. This enables the generalization of the sound-field production to a rigid surface with an arbitrary form factor and microphone array size.
@@
The paper is organized as follows. In section II, we lay down the theoretical foundation of the work. The details of the proposed framework are described in section III. Then, we present the validation results in section IV. Finally, we describe few engineering applications in section V. The following notations are used throughout the paper. A bold lower-case letter denotes a column vector, while a bold upper-case letter denotes a matrix. $M$ always refers to the number of microphones. The independent variables $t$ and $\omega$ refer to time and frequency respectively. Additional notations are introduced when needed.
@@ II-A Acoustic Plane Waves
Acoustic plane waves are eigenfunctions of the homogenous Helmholtz equation. Hence, they constitute a powerful tool for analyzing the wave equation. Further, a plane wave is a good approximation of the wave-field emanating from a far-field point source [8]. The acoustic pressure of a plane wave with vector wave number $\bf{k}(\theta,\phi)$ (where $\theta$ and $\phi$ correspond respectively to polar azimuth and elevation of the direction of propagation) is defined at a point ${\bf{r}}=(x,y,z)$ in the three dimensional space as [9]:
$\psi({\bf{\omega,\theta,\phi,r}})\triangleq p_{0}(\omega)e^{-j{\bf{k}}^{T}{\bf{r}}}$ (1)
where $p_{0}(\omega)$ is a real-valued frequency dependent scaling. The plane wave decomposition has been used for approximating point-source seismic recording [10, 11, 12], and sound field reproduction [8, 13, 14, 15]. A local solution to the homogenous Helmholtz equation can be approximated by a linear superposition of plane waves of different angles of the form [11, 16]:
$p(\omega,{\bf{r}})=\sum_{l\in\Lambda}\alpha_{l}(\omega)\ \psi\left(\omega,\theta_{l},\phi_{l},\bf{r}\right)$ (2)
where $\Lambda$ is a set of indices that defines the directions of plane waves $\{\theta_{l},\phi_{l}\}$, each $\psi(.)$ is a plane wave as in (1), and $\{\alpha_{l}\}$ are complex-valued scaling factors. We will refer to the wave-field in (2) as the free-field acoustic pressure. The decision variables in this approximation are $\left\{\Lambda,\{\alpha_{l}\}_{l\in\Lambda}\right\}$.
II-B Device Acoustic Dictionary
Generalizing the free-field plane wave expansion in (2) to include the scattering due to the device surface, requires computing the device acoustic response to each plane wave. The device response to all plane waves in the three-dimensional space is collectively referred to as the device acoustic dictionary. The total wave-field at any point on the device surface when the device is impinged by an incident plane wave $\psi(\omega,\theta,\phi,{\bf{r}})$ has the general form:
$p_{t}(\omega,\theta,\phi,{\bf{r}})=\psi(\omega,\theta,\phi,{\bf{r}})+p_{s}(\omega,\theta,\phi,{\bf{r}})$ (3)
where $p_{t}$ and $p_{s}$ refer to the total and scattered wave-field respectively. $p_{t}$ can be computed numerically by inserting (3) in the Helmholtz equation and solving for $p_{s}$ with appropriate boundary conditions. If a microphone array of size $M$ is mounted on the device surface, and the microphone port size is much smaller than the wavelength, then each microphone can be approximated by a point on the device surface. In this case, the total field, ${\mathbf{p}}_{t}(\omega,\theta,\phi)$, at the microphone array, due to an incident plane wave $\psi(\omega,\theta,\phi,{\bf{r}})$, is a vector of size $M$ whose entries are the corresponding total field at the coordinate values, $\bf{r}$, of each individual microphone. The device acoustic dictionary of a device is composed of vectors of total acoustic wave-field. The device acoustic dictionary is computed using numerical acoustic simulation with Finite Element Method (FEM) or Boundary Element Method (BEM) with device CAD to specify the device surface. The details and validation results are described in [3].
@@ II-B Device Acoustic Dictionary
The output of the above model is the plane wave dictionary of the device
@@
where each entry in the dictionary is a vector of length $M$, and each element in the vector is the total acoustic pressure at one microphone in the microphone array when a plane wave with ${\bf{k}}(\theta_{l},\phi_{l})$ hits the device. The dictionary also covers all frequencies of interest typically up to $8$ kHz. Note that, the acoustic dictionary can accommodate any form factor of the surface, and any geometry of the microphone array as it utilizes acoustic simulation with the CAD of the device surface, and the coordinates of the microphone array.
III Proposed Framework
III-A Overview
The proposed sound synthesis methodology generalizes the plane wave expansion in (2), which summarizes room acoustics at a point in a room, to include the impact of scattering due to the device surface. This generalization yields the combined acoustic effect of the room and the scattering on the device surface. If the device dimensions are much smaller than the room dimensions, then secondary reflections due to the device surface are negligible, and the device impact on the room acoustics could be ignored. Hence, even after introducing the device into the room, the directions and weights of the free-field acoustic plane waves in (2) do not change. However, because of the device surface, each plane wave $\psi\left(\omega,\theta_{l},\phi_{l},{\bf{r}}\right)$ in (2), has an acoustic fingerprint at the device microphone array, which is the corresponding entry in the device dictionary, $\boldsymbol{\beta}_{l}(\omega)$. Hence, if the device is placed at a point in the room whose sound-field is expressed as in (2), then the observed sound field vector at the device microphone array is
@@ III-A Overview
The transition from the sound-field in (2) in the absence of the device to the sound field in (5) in the presence of the device is the technical foundation of the proposed synthesis framework. This transition is enabled by the linearity of the wave equation and the introduction of the device acoustic dictionary as described in section II-B.
Note that, if another device with acoustic dictionary $\mathcal{D}^{(2)}\triangleq\{\boldsymbol{\beta}_{l}^{(2)}(\omega):\forall\ \omega,l\}$ is placed at the same point in the room, then the observed sound-field at the second microphone array can be expressed as
@@
where only the mapping through device dictionary changes, while the directions, $\Lambda$, and weights, $\{\alpha_{l}\}$, of constituent plane waves do not change. This is the essence of the proposed methodology that separates room and device components. Hence, the three steps for sound-field synthesis at a microphone array mounted on a device that is placed at a point in the room are as follows:
@@
    Compute the free-field plane wave expansion at a point as in (2). This summarizes room acoustics at a point in the room. A procedure for computing this expansion is described in section III-B. This process is repeated for every new point in a room, and it is independent of the device under test.
@@
    Generate the acoustic dictionary of the device under test as described in section II-B. This is computed once per device and it is independent of the room.
@@
    For each room position, combine the plane wave expansion with the device dictionary as in (5) to synthesize the sound-field at the device microphone array.
Repeating step $1$ above for multiple rooms and multiple positions within each room generates a room database. This database is generated only once, then it could be reused in evaluating and generating data for new devices.
III-B Acoustic Plane Wave Decomposition
The main technical hurdle in the proposed framework is computing the plane wave expansion (2) at a point in a room with a source signal emanating from another point in the room. In the proposed framework, this is computed through a data capture using a large microphone array of $32$ microphones mounted on a sphere (EigenMike) [21]. The large microphone array is necessary to mitigate the creation of an underdetermined system of equations in recovering the constituent plane waves. It was found experimentally that $20$ to $30$ plane waves are sufficient for an accurate approximation of the sound field (with reconstruction error less than $-30$ dB for frequencies up to $8$ kHz). The plane wave decomposition problem is formulated as an optimization problem whose objective is minimizing the difference in the least square sense between observed and synthesized sound fields. If the observed sound field of the EigenMike at frequency $\omega$, is ${\mathbf{y}}(\omega)$, then the objective function has the form
@@ III-B Acoustic Plane Wave Decomposition
where $\left\{\bar{\boldsymbol{\beta}}_{l}(\omega)\right\}$ are the entries of the EigenMike acoustic dictionary at frequency $\omega$. The decision variables are the set of indices $\Lambda$, and the corresponding weights $\{\alpha_{l}\}$. The L1-regularization in (7) is added to stimulate a sparse solution as $|\Lambda|<30$ is much smaller than the dictionary size, which is in the order of $10^{3}$. The objective function can be put in matrix form as:
@@
III-C Room Impulse Response (RIR) Computation
RIR aims at modeling the acoustic channel between source and receiver as a linear time-invariant system. The RIR combines both room acoustics and scattering due to device surface, and it is computed once for a given device and a given source/receiver positions in a room. It is a multichannel transfer function where the number of channels equals the size of the microphone array. For RIR computation, a special source signal that covers the whole frequency spectrum, e.g., white noise or Golay sequence [27], is utilized. For a source signal $x(t,\omega)$, the EigenMike is utilized to generate the time-frequency map of the plane wave decomposition as described in the previous section. For a device under test, this time-frequency map is combined with the device dictionary to generate the multichannel output signal ${\mathbf{y}}(t,\omega)$ as in (5). The transfer function between $x(t,\omega)$ and ${\mathbf{y}}(t,\omega)$ is computed using system identification techniques. For example, by applying Wiener-Hopf equation in the frequency domain [28], we get